Q1) The relationship between sleep hours and overall happiness level on the next day was estimated as below in a linear format.

(Happiness level is measured through a five-point scale: 1 = Extremely unhappy, 5 = Extremely happy)

Overall happiness level = 0.27 * Sleep hours + 2.34

(The p-value for the coefficient of Sleep hours is 0.02.)

How can the relationship between overall happiness level and sleep hours be interpreted?

A) As a person sleeps 1 more hour, his or her happiness level is expected to increase by 2.34 points.

B) There is no meaningful relationship between happiness level and sleep hours from a statistical point of view.

C) As a person sleeps 1 more hour, his or her happiness level is expected to decrease by 0.27 points.

D) As a person sleeps 1 more hour, his or her happiness level is expected to increase by 0.27 points.

Q2)

The relationship between sleep hours and overall happiness level on the next day was estimated as below in a linear format.

(Happiness level is measured through a five-point scale: 1 = Extremely unhappy, 5 = Extremely happy)

Overall happiness level = 0.27 * Sleep hours + 2.34

(The p-value for the coefficient of Sleep hours is 0.02.)

Based on this equation, what is the expected happiness level of a person on a certain day, when he slept 7 hours the day before?

A) 4.17

B) 3.89

C) 4.23

D) 2.34

Q3

The relationship between annual salary (in US dollars) and gender was estimated as below in a linear format.

(Female was coded as 0 and Male was coded as 1.)

Annual Salary = 23,472 * Male + 44,728

(The p-value for the coefficient of Male is 0.24.)

A) How can the relationship between gender and annual salary be interpreted?

B) Male employees are expected to receive 44,728 dollars more than female employees

C) Male employees are expected to receive 23,472 dollars less than female employees

D) Male employees are expected to receive 23,472 dollars more than female employees

There is no meaningful difference in annual salary between male and female employees from a statistical point of view.
Answer #1

Q1)
Overall happiness level = 0.27* Sleep hours + 2.34

Ans: D) As a person sleeps 1 more hour, his or her happiness level increases by 0.27 points

The slope or the coefficient of independent variables gives the change in dependant variable(overall happiness) for unit change in independent variable.

Q2)
Overall happiness level = 0.27*7+2.34 = 4.23
Ans: C)4.23

Q3)
When code = 0 we get the salary of female
Annual Salary (female) = 23472*0+44728 = 44728

When code =1, we get the salary of male
Annual salary(male) = 23472*1+44728 = 68200

Ans:D) Male employees are expected to receive 23472 dollars more than female employees